OpenStudy (anonymous):

cube root of 389017 by estimated method(with steps

OpenStudy (anonymous):

First divide the number into groups of three digits beginning from the ones' digit as follows : 389 | 017 Group II | Group II

OpenStudy (anonymous):

Group I ends in 7. Now we have to find a digit A such that A*A*A ends in 7. With a little hit and trial we find 3*3*3 = 27 ends in 7.

OpenStudy (anonymous):

Group II is 389. Now we find a number B such that B*B*B ≤ 389 With a little hit and trial we find 7*7*7 = 343 ≤ 389 8*8*8 = 512 which is more than 389

OpenStudy (anonymous):

From Group I we got the digit 3 and we take it as ones' digit From Group II we get the digit 7 and we take it as tens' digit Now we calculate 73^3 and find it is 389017 so cube root of 389017 is 73.

OpenStudy (anonymous):

Another way could be 389,017 < 400,000 50*50*50 = 125,000 60*60*60 = 216,000 70*70*70 = 343,000 80*80*80 = 512,000 so cube root of 389017 lies between 70 and 80 now the no 389017 ends in 7 so we find by hit and trial (as above) that 3*3*3 = 27 so we try 73^3 and find it is equal to 389017 so cube root of 389017 is 73

OpenStudy (ybarrap):

This follows Harkirat's initial logic and is documented here: http://mbawrites.com/mba/cat/how-to-calculate-cube-roots-in-less-than-5-seconds/ I summarize the process for this case: Memorize cubes up to nine and find the cube that is less then 389; this would be 7, since 7^3 is 343 and 8^3 is 512. So 7 is the ten's digit Memorize the last digit of the cubes from 1 to 9, this becomes the unit portion of the answer Since the last digit is 7, the units digit is 3. Ans. 73